OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5). −6 4.5 6 14

OpenStudy (anonymous):

14

OpenStudy (aroub):

-2(-5)+4 -10+4 -6

OpenStudy (aroub):

-6

OpenStudy (anonymous):

aroub enter doesnt work

OpenStudy (anonymous):

hehe

OpenStudy (anonymous):

so medal given

OpenStudy (anonymous):

It's 14 bro

OpenStudy (aroub):

lol yea!! :( :( no well actually its 14 sorryy

OpenStudy (anonymous):

is it i dont doubt aroub

OpenStudy (anonymous):

so its is just plugged it in

OpenStudy (aroub):

lol!

OpenStudy (anonymous):

see here: we have given the value of x is - -5 . plug in the value of x in the equation -2x+4 = -2(-5) + 5 = 10 + 5 = 14

OpenStudy (anonymous):

got it